OpenStudy (anonymous):

(algebra) Consider two numbers 24/x and 24/x+1 , where x is a positive integer. If the difference between the numbers is 1/3. A) show that x²+x -72 = 0 B) Hence find the value of X, where X=0 and X is a positive integer.

OpenStudy (anonymous):

if the difference of the two numbers is \(\frac{1}{3}\) you can write \[\frac{24}{x}=\frac{24}{x+1}+\frac{1}{3}\]

OpenStudy (anonymous):

how do u do the 24/x thing?

OpenStudy (anonymous):

thx btw *

OpenStudy (anonymous):

\frac{a}{b}